Four houses on one side of a street have addresses that are consecutive odd numbers. Find each address if the sum of these addresses is 23,672.
Thank YOU :)

Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12569)

Let x, x+2, x+4 & x+6 be the 4 numbers.
x+x+2+x+4+x+6=23,672
4x+12=23,672
4x=23,672-12
4x=23,660
x=23,660/4
x=5,915 is the lowest house number.
5,915+2=5,917 is the second house number.
5,915+4=5,919 is the third house number.
5,915+6=5,921 is the largest house number.
Proof:
5,915+5,917+5,919+5,921=23,672
23,672=23,672

Answer by Alan3354(30993)

23672/4 = 5918 which is the center of the 4 numbers
--> 5915, 5917, 5919, 5921